OpenStudy (anonymous):

Which second-degree polynomial function has a leading coefficient of negative 1 and root 5 with multiplicity 2? A. f(x)=-x^2+10x+25 B. f(x)=-x^2+10x-25 C. f(x)=-x^2-10x+25 D. f(x)=-x^2-10x-25

OpenStudy (rational):

Look at option B : \[\large f(x) = -x^2 + 10x - 25 = -(x^2-10x+25) = -(x-5)^2\]

OpenStudy (anonymous):

Thanks

OpenStudy (rational):

Since the exponent on factor \(x-5\) is \(2\), the multiplicity of root \(5\) is \(2\)